answers. If the answer is a number, do not round. (a) T. +5 dr = (b) est dir = (o (1 + 2) dt = dr
Answer #1

Sall T- नि se dae M+5 Substituting 245=t → २ = dt मन dt 2 I - dt 2 । -४ fctjedt - 와 (t) Y2 + C. Scanned with CamScanner Jt+cI Ja²+5+c (:: += x++) (6) 3 54 I = Se e dae -2 3 50 e ما به -2 3 50 s ces u -2 H = I Ś[es-elo] Ans. (c) settajat CS Scanned wNow, By using Leibnitz Differentiation under Rule of entegral sign ule get- (éta ). d (o?)-(1+2). (1) dae dae = (p+8). (ax) –